OpenStudy (anonymous):

A triangular plot of land has boundary lines 45 meters, 60 meters, and 70 meters long. The 60 meter boundary line runs north-south.Is there a boundary line for the property that runs east-west?

OpenStudy (amistre64):

is it a 90?? 60^2 + 45^2 = 70^2 or maybe 60^2 + 70^2 = 45^2

OpenStudy (amistre64):

0 + 5 = 0...first one aint gonna happen

OpenStudy (amistre64):

0+0 = 5....aint gonna happen here either chances are NO!!